# Exploring the Nostr Ecosystem: A Study of Decentralization and Resilience

### v1

@@ 6.2. Evaluating Strategic Replication Reduction
AS-Based Relay Selection.  We next test the strategy where we bias the selection of replicas based on the AS that hosts the relay. Intuitively, we would want a post to be replicated across different ASes to prevent a central points of failure. Consequently, rather than selecting randomly, for every $P\rightarrow S$ with $|S|>N$, we construct subset $S^{\prime}$ such that relays in $S^{\prime}$ span as many ASes as possible, and let $P\rightarrow S^{\prime}$. To do this, we start by selecting one random relay from $S$ for each possible AS. If we do not have enough, we randomly select additional relays from $S$ until the total number equals N. Utilizing this approach, we once again run the experiments in 5 scenarios with $N$ equal to 2, 5, 10, 15, and 20, respectively.
The availability results are depicted in Figure 9b. It is evident that the outcomes are significantly superior to random selection, and for $N\geq 10$, they closely resemble the original results. To quantify this, we find $D^{A}_{relay}(10,50)$ is -0.39%, and $D^{A}_{AS}(10,30)$ is -0.53%, while for random relay selection, $D^{A}_{relay}(10,50)$ is -1.15%, and $D^{A}_{AS}(10,30)$ is -8.38%. These figures confirm that the reduction in availability is minimal. These results suggest that by choosing a strategy that prioritizes post availability, it is possible to choose smaller values of $N$ without negatively impacting post availability.
Follower-Based Relay Selection.  Sometimes, the priority is not the availability of the post, but rather, ensuring that the user’s followers are able to view it. Recall, Nostr requires some degree of coordination in which publishers push posts to the same relays that consumers request them from. To address this, we explore another strategy, in which we bias replication towards relays where followers reside.
First, for each user, we create a relay list $L_{U}$, ranking it based on the number of the user’s followers using that relay (we define that a user is using a relay if they have at least one post on the relay). Then, for every $P\rightarrow S$ where $|S|>N$, we construct a subset, $S^{\prime}$, by selecting relays from $S$ based on their priority as indicated by $L_{U}$, where $U$ is the author of $P$. Utilizing this strategy, we run the experiments again in 5 scenarios with $N$ equal to 2, 10, 15, 20, and 25, respectively. The availability results are depicted in Figure 9c.
We observe that the post availability using this relay selection strategy is less optimal. The values of $D^{A}_{relay}(20,50)=-5\%$ and $D^{A}_{AS}(20,30)=-17.74\%$ indicate that the availability is lower than that of the original. Furthermore, this outcome is also less optimal compared to the random relay selection approach, where $D^{A}_{relay}(20,50)=-0.05\%$ and $D^{A}_{AS}(20,30)=-2.65\%$. This outcome seems intuitive, as this strategy tends to centralize replications on a few popular relays. Thus, this confirms that certain relay selection strategies can result in availability that is considerably worse than random.
Evaluation Summary.  Overall, the evaluation demonstrates that strategic replication is a practical approach to reduce post replication redundancy, without negatively impacting availability. Specifically, for our first evaluation objective, the results confirm that when $N=20$, post availability is not significantly impacted. For random selection, the decrease of availability is only 2.65%. Furthermore, employing a relay selection algorithm that prioritizes availability allows us to obtain less than 0.01% decrease in availability and thus opt for smaller $N$ values. For our second evaluation objective, our results suggest that the best selection strategy is AS-based with $N=20$. In this case, 380 million post replications can be eliminated with little negative impact to the system. That said, we hope that others will continue to explore alternative selection strategies.
6.3. Improving Retrieval Traffic Waste with Optimistic Retrieval
To tackle the issue of duplicated retrieval of posts by clients (as demonstrated in §5.4) we propose what we refer to as optimistic retrieval. In the vanilla Nostr implementation, user clients will retrieve the posts of every user they follow from all relays in their relay list (outlined in the pseudo code in Algorithm 4). This creates substantial communications overhead for both clients and relays. To address this, we propose an optimistic retrieval method, whereby the user retrieves posts from a (potentially different) set of $N$ relays for each of their following users respectively. By lowering $N$ we can reduce the wasted traffic, although naturally it is necessary to intelligently select which relays to check. For instance, imagine if user $A$ follows users $B$ and $C$; and $A$’s relay list includes three relays ($R_{1}$, $R_{2}$, and $R_{3}$). $A$’s client may choose to retrieve $B$’s posts only from $R_{1}$ and $R_{2}$, and $C$’s posts only from $R_{2}$ and $R_{3}$, instead of retrieving $A$ and $B$’s posts from all three relays. Clearly, if $N$ is small, the amount of wasted traffic can be significantly reduced.
This method consists of three parts. The first is defining a new type of tag [^2] https://github.com/nostr-protocol/nips?tab=readme-ov-file#standardized-tags in the Nostr event object, which we call priority. The second part is about how to select the $N$ relays. The third part is the actual optimistic retrieval procedure executed by the clients.
The Priority Tag.  If clients do not connect to all relays, it is possible that they may (theoretically) miss certain posts. This can be damaging for extremely important or delay-intolerant content To mitigate this, we propose a new tag which we add to the Nostr event object (recall, there are many event types in Nostr that can be stored in relays). This boolean tag allows the relay and client to distinguish important events in the filter within the subscription. In practical terms, a normal event is published without a priority tag. If a user deems an event to be important, they will publish it with priority tag. For these high-priority events, retrieval client will not employ optimistic retrieval but will still retrieve it from all relays in the relay list.
Relay Selection.  To help others select the $N$ relays for retrieval, each user will publish (and update) an event containing a short list of ”recommended relays”. This is pushed to a relay like any other event. This is already specified as a kind-10002 event (Nostr, 2023e), and there is therefore no need to for us to introduce additional functionality. Other clients (the followers of this user) will then parse this information and utilize the list to construct the set of $N$ relays for optimistic retrieving this user’s posts. This event should be published with a priority tag to ensure it is not overlooked.
Optimistic Retrieval Procedure.  We detail the full pseudo code in Algorithm 5. Based on the aforementioned list of ”recommended relays”, the retrieval client first construct a set $S_{U}^{R}$ of $N$ relays for each following user $U$ respectively. For each relay $R$ in the retrieval client’s relay list, it only adds a subscription of ”low-priority” events (without the priority tag) for a following user $U$ if $R\in S_{U}^{R}$. On the other hand, for each relay $R$, it still adds a subscription of ”high-priority” events (with the priority tag) for all following users.
6.4. Evaluating Optimistic Retrieval
By adopting optimistic retrieval, we can reduce wasted traffic. However, there is a key trade-off in terms of the completeness of the retrieved posts. This is because the user only retrieves posts from a subset of the relays, rather than from all relays in their list. As a result, they may miss some posts (though deemed to be ”low-priority”) that are not hosted by this subset of relays.
Therefore, to evaluate the optimistic retrieval method, we concentrate on two aspects. The first is the completeness of retrieved post set, i.e., the proportion of posts successfully retrieved using optimistic retrieval. Our aim is to demonstrate that, even with a small number of $N$ relays, optimistic retrieval can still download the vast majority of posts that the user follows. The second aspect measures the traffic savings resulting from the optimistic retrieval.
Completeness of Retrieved Posts.  We first assess whether it is possible to retrieve a high percentage of available posts of each user, through optimistic retrieval. In order to achieve this, we must demonstrate that it is possible to retrieve a high percentage of posts for each user (post publisher) respectively from their ”recommended relays”.
Thus, we start by examining the proportion of posts that a set of relays can cover for a user. We denote $S_{U}^{P}$ as the set of all posts of the user $U$. We denote $S_{R}^{P}$ as the set of all posts contained within the relay $R$. The intersection, $S_{U}^{P}\cap S_{R}^{P}$ represents the posts made by user $U$ that are covered by relay $R$. Then, we define the coverage of $R$ for $U$ as $\frac{|S_{U}^{P}\cap S_{R}^{P}|}{|S_{U}^{P}|}\times 100\%$, that is the percentage of the posts of $U$ that is covered by $R$. Similarly, for a set of relays $S^{R}$, we first compute the set of all posts contained within the relays in $S^{R}$, as $S_{S^{R}}^{P}=\bigcup_{R\in S^{R}}S_{R}^{P}$. Then the coverage of $S^{R}$ for $U$ is $\frac{|S_{S^{R}}^{P}\cap S_{R}^{P}|}{|S_{U}^{P}|}\times 100\%$.
First, for each user $V$, we evaluate the coverage of $V$’s recommended relays for $V$. This corresponds to the proportion of posts that $V$’s follower can retrieve using optimistic retrieval. We assume the user employs a naive method that they recommend the relays with largest number of their posts. For each user, we construct 5 sets with the best one, two, three, four and five relays respectively. Afterwards, we create a plot for the CDF of the coverage offered by such five sets of relays for each user. The result is in Figure 10a. Here, we observe that, With the top four or top five relays, nearly 90% of users can attain complete 100% coverage. The results confirm that by strategically opting for a small number of relays, the post coverage for the majority of users can be satisfactory.
Then, we construct the follower graph $G$, and evaluate the post retrieval completeness over $G$. We evaluate the completeness of each following edge $(U,V)$. Specifically, let $S^{P}_{V}$ denote the set of all posts by $V$, and $O^{P}_{V}$ denote the set of $V$’s posts that can be retrieved by $U$ using optimistic retrieval. The completeness is then computed as $\frac{|O^{P}_{V}|}{|S^{P}_{V}|}\times 100\%$. Figure 10b shows the CDF of completeness of the edges. We observe similar result that nearly 90% of following edges achieve 100% coverage.
The results verifies that optimistic retrieval can achieve high completeness with a small number of $N$ ($N=5$). Moreover, relay selection strategy for ”recommended relays” can be as simple as selecting the top $N$ relays that have the highest number of posts by the user. Overall, optimistic retrieval is a feasible and practical method for enhancement.
Saved Traffic.  Now we measure the traffic savings that our optimistic retrieval achieves. To do this, we use the same approach as described in §5.4, while assuming optimistic retrieval is employed and $X$of posts are of ”high-priority”. Specifically, in the original method in §5.4, for a following edge $(U,V)$, we retrieve from $S_{U}^{R}$, which represents the set of relays utilized by user $U$. Now, for ”high-priority” event, we still use $S_{U}^{R}$. While for normal posts, we use $S_{V_{N}}^{R}$, which represents the recommended $N$ relays for user $V$ (top $N$ relays with largest number of $V$’s posts).
Table 1 presents the results for $N=$ 2, 3, 4, and 5, assuming 5% of posts are with priority tag (see Appendix A.4 for the tables for other percentages of priority tag). The table clearly illustrates that even when $N$ is set to 5, implementing optimistic retrieval has the potential to decrease the duplicated retrieval count by 88.3%, while still achieving 99.4% of post completeness. This confirms the substantial improvement that can be achieved through optimistic retrieval, without compromising the completeness of the retrieved posts. This is important, particularly as many relays are lightweight and may have to pay for metered bandwidth.
@@ 6.4. Evaluating Optimistic Retrieval
Wasted Count Wasted Size
Overall Completeness (%) Wasted (Billion) % of Original Wasted (TiB) % of Original
Original - 218.3 - 143.9 -
$N=2$ 99.0 14.5 6.6 9.6 6.7
$N=3$ 99.2 18.2 8.3 12.0 8.3
$N=4$ 99.3 21.9 10.0 14.5 10.1
$N=5$ 99.4 25.5 11.7 16.9 11.7
Table 1. The post completeness and potentially wasted count and size of posts over the following graph, when adopting optimistic retrieval with N = 2, 3, 4, and 5, with 5% ”high-priority” posts.
@@ 7. Related Work
To the best of our knowledge, this research is the first large-scale analysis of the Nostr ecosystem. However, efforts to establish decentralized online social networks date back to the inception of platforms like Safebook (Cutillo et al., 2009), PeerSoN (Buchegger et al., 2009), LotusNet (Aiello and Ruffo, 2012), and LifeSocial.KOM (Graffi et al., 2011), which were based on peer-to-peer (P2P) technologies. Since, federated networks such as Mastodon and PeerTube have risen, forming what is known as the Fediverse (Mansoux and Abbing, 2020). These utilize the W3C ActivityPub protocol (W3C, 2018) to facilitate server-to-server communication.
Measurements of the Fediverse.  Several recent works have explored these emerging federated social networks. Closest to our work is an initial investigation highlighting the centralization tendencies within Mastodon (Raman et al., 2019), with a focus on structural and robustness aspects. Growth patterns of Mastodon in its initial stages were the subject of another study (Cava et al., 2021). Investigations into user behaviors across different instances have also been conducted (Cava et al., 2022a, b). The aspect of content moderation has been previously examined in the context of Pleroma, another service within. Hassan et al. shed light on new challenges (Hassan et al., 2021) and suggest a preliminary solution, while Zia et al. (Zia et al., 2022) offer a model for shared moderation efforts.
Nostr differs from other Fediverse applications in several aspects. Most notably, relays in Nostr do not communicate with each other. However, users have the ability to communicate directly with multiple relays and switch between them at no cost. This distinctive feature of Nostr results in user accounts that are not dependent on any specific relay or centralized entity. It also grants users the freedom to send and replicate posts across various relays, thereby enhancing availability and resistance to censorship.
P2P Social Networks.  Meanwhile, P2P decentralized social networks such as Secure Scuttlebutt (SSB) still exists, offering a P2P event-sharing protocol and an architecture for social applications (Tarr et al., 2019). There are also blockchain-based social networks, leveraging cryptocurrency to incentivize users to share content and to prevent malicious behaviors. Nostr is rather different, and does not introduce any blockchain solutions (although users can use cyrptocurrency to pay for services). Research has been conducted on various on-chain social networking applications, including memo.cash (Zuo et al., 2023), Steemit (Guidi et al., 2020), and Sapien (Jaatun et al., 2022). Additionally, to address privacy concerns within native blockchains, there have been recent proposals to integrate blockchain with social media applications (Arquam et al., 2021; Guidi, 2020; Jiang and Zhang, 2019). While Nostr employs public-private key encryption to sign posts for verifiability, it diverges from utilizing P2P technology or blockchain technology. Unlike P2P social networks such as SSB, which are more intricate to set up and demand each server/user to store the entire chain of posts for validation of new ones, Nostr is simpler in concept, lightweight, and easier to use.
8. Conclusion
This paper has presented a comprehensive measurement study of the Nostr ecosystem, an emerging and innovative decentralized social network. Our focus has centered on investigating two key innovations introduced by Nostr. First, we investigated the role of independent Nostr relays distributed across various countries. Our study of Nostr relays has revealed significant gains in decentralization, with a relatively even distribution of posts and users across relays. Similarly, we have observed a relatively even distribution of relays across countries and autonomous systems. However, the challenge of unstable relay availability persists, with 20% experiencing substantial downtime. We conjecture that this might be attributed to financial sustainability issues especially for free-to-use relays. Indeed, we find that 95% of such relays would struggle to cover operational costs from the publication donations.
Second, we studied the replication of posts across multiple relays, a distinct feature of Nostr, along with its associated tradeoffs concerning post availability and overhead. We found that Nostr’s unique post replication approach does enhance availability, but introduces significant overhead. Our findings indicate that post availability surpasses traditional Fediverse applications, even during extensive relay or AS outages. However, the substantial replication overhead (average of 34.6 relays per post) leads to wasteful storage and bandwidth consumption. Addressing this concern, we proposed incorporating innovative features in post publication. Our evaluation has shown that these enhancements are effective, resulting in a remarkable 380 million reduction in post replications, while sustaining the same level of availability. Additionally, we have recommended a further refinement in post retrieval. This significantly decreases wasteful retrievals from 98.2% to 11.7% based on our evaluation. These proposed improvements underscore the considerable potential for optimizing Nostr’s performance.
This paper constitutes the first large-scale study of Nostr. Consequently, there are a number of lines of remaining future work. Most notably, we wish to pilot and deploy our system innovations at scale, to understand their impact on things like storage redundancy. We also note that a number of secondary applications have been built a top of Nostr, e.g., chat rooms and live streaming. Investigating their usage form a core pillar of our future work.

### v2

@@ 6.1. Improving Storage Redundancy
We experiment with two strategies for selecting $S^{\prime}$: random (as a baseline) and our proposed AS-based strategy. Specifically, with the random strategy, for every $P\rightarrow S$ with $|S|>N$, we construct subset $S^{\prime}$ by randomly selecting $N$ relays from $S$, and let $P\rightarrow S^{\prime}$. For the the AS-based strategy, for every $P\rightarrow S$ with $|S|>N$, we construct subset $S^{\prime}$ such that the relays in $S^{\prime}$ span as many ASes as possible, and let $P\rightarrow S^{\prime}$. To do this, we start by selecting one random relay from $S$ for each possible AS. If we do not have enough, we randomly select additional relays from $S$ until the total number equals N. This is designed to be inherently simple and not require coordination across multiple clients or relays.
Figure 9. Percentage of available posts when top X relays are removed and when top X ASes are removed, if the maximum number of post replications is (a) reduced to N randomly; (b) reduced to N based on AS.
Evaluation Results.  We use the same method as in §5.2 to simulate relay failures and thus compare the post availability under attack to the original. Figure 9 shows the results of the impact on availability when removing either (i) individual relays; or (ii) entire ASes (and therefore all relays hosted within). To quantify the difference in availability between the new scenario and the original, we begin by defining $A(N,X)$ as the availability when the maximum replication is restricted to $N$, and the top $X$ relays are removed. This corresponds to each point depicted in the figure. Subsequently, to capture the difference in availability between the given scenario and the original, we introduce $\Delta^{A}(N,X)=A(N,X)-A(O,X)$, where $O$ denotes the original scenario. In order to encapsulate the overall difference between a line in the figure and the original, we define $D^{A}_{relay}(N,X)=\Sigma^{X}_{i=1}\Delta^{A}(N,i)/X$. This captures the average difference in availability across all $X$ values. Similar to $D^{A}_{relay}$, we also define $D^{A}_{AS}(N,X)$ for the removal of $X$ ASes.
Figure 9b plots the results for the AS-based strategy. Even when removing several relays or ASes, if the maximum number of post replicas is limited to $N\geq 10$, the post availability (percentage of available posts) is very close to the original. That is, throughout the measured $X$ values, $D^{A}_{relay}(N,X)<1\%$ and $D^{A}_{AS}(N,X)<1\%$ for $N\geq 10$. Additionally, this is also significantly better than the random strategy, as depicted in Figure 9a, especially in the scenario where several ASes (and all relays hosted within) are removed. To quantify this, we find $D^{A}_{relay}(10,50)$ is -0.39%, and $D^{A}_{AS}(10,30)$ is -0.53%. In contrast, for random relay selection, $D^{A}_{relay}(10,50)$ is -1.15%, and $D^{A}_{AS}(10,30)$ is -8.38%. This outcome is anticipated, as this relay selection strategy directly opposes scenarios where the entire AS is removed.
Evaluation Summary.  The results confirm that our proposed solution is effective. For reasonable values of $N$ (such as 15 and 20), we dramatically reduce the post redundancy while maintaining high levels of post availability and censorship resistance against our simulated relay-level and AS-level failures. Further, by choosing our proposed AS strategy, it is possible to choose smaller values of $N$ such as 10. In the case where the number of replications is restricted to $N=20$, 380 million (61.7%) post replications can be eliminated.
However, it is important to note that this AS-spanning replication optimization could introduce a load imbalance, particularly for relays that are the sole one hosted in their AS. This could open up potential threats. For example, an adversary could exploit this by identifying and targeting these unique relays in small ASes to disproportionately impact the overall network. Our current resilience evaluation assumes whole AS-level failures and does not model this more surgical targeted relay attack. Therefore, Nostr’s resilience under more targeted adversarial strategies remains an open question.
6.2. Improving Retrieval Traffic Waste
The issue of duplicated retrieval of posts by clients (§5.4) arises from the fact that, in the vanilla Nostr implementation, clients retrieve the posts of every user they follow from all relays in their relay list. This creates substantial communications overhead for both clients and relays. An intuitive solution might be for a client to send a filter requesting the relay to exclude any prior posts they already possess. However, in practice, such a solution is impractical. This is because clients connect in parallel to all relays, to enable real-time communications. Sending a filter would require a client to connect to each relay sequentially one-by-one (so that the filter can be updated). This could result in a delay of several minutes for the client to refresh the timeline, which many users may find unacceptable. It also introduces additional complexity and network load (e.g., sending a filter also cost additional bandwidth) on both the client and relay-side, which Nostr strives to avoid.
Thus, we propose a design improvement that avoids any additional client-relay communication, where the client only retrieves posts from only a small subset of relays (rather than from all relays in their list). This, however, demands the intelligent selection of which relays to contact. To select which $N$ relays to retrieve posts of user $V$ from, we first rank all relays based on the number of user $V$’s posts stored there. We then select the top $N$ relays that host the most posts by user $V$. We posit that this lightweight strategy increases the likelihood of accessing most posts. However, even with the best relays, posts can be missed. Thus, to evaluate this trade-off, we test our proposed design by simulating the scenario where clients only retrieve posts from a selected subset of $N$ relays.
@@ 6.2. Improving Retrieval Traffic Waste
Evaluation of Post Coverage.  We first evaluate the strategy from the post sender’s perspective, where we examine the proportion of posts that our approach of selecting the $N$ relays can cover for a sender. We denote $S_{U}^{P}$ as the set of all posts of user (sender) $U$. We denote $S_{R}^{P}$ as the set of all posts contained within the relay $R$. The intersection, $S_{U}^{P}\cap S_{R}^{P}$ represents the posts made by sender $U$ that are stored by relay $R$. Then, we define the coverage of $R$ for $U$ as $({|S_{U}^{P}\cap S_{R}^{P}|}\ /\ {|S_{U}^{P}|})\times 100\%$, i.e., the percentage of the posts of $U$ that is covered by $R$. Similarly, for a set of relays $S^{R}$, we first compute the set of all posts contained within the relays in $S^{R}$, as $S_{S^{R}}^{P}=\bigcup_{R\in S^{R}}S_{R}^{P}$. Then the coverage of $S^{R}$ for $U$ is $({|S_{S^{R}}^{P}\cap S_{R}^{P}|}\ /\ {|S_{U}^{P}|})\times 100\%$.
For each user (sender) $V$, we evaluate the coverage of the selected $N$ relays (the top $N$ relays ranked by the number of $V$’s posts) for $V$. We construct 5 sets with the $N=$ 1, 2, 3, 4, and 5, respectively. Figure 10a presents the CDF of the coverage offered by the five set sizes of relays for each sender. We see that, by retrieving solely from the top four or top five relays, 90% of senders can ensure that all recipients see their posts (i.e., 100% coverage). This confirms that by strategically opting for a small number of relays to push posts to, the post coverage for the majority of senders is satisfactory.
comple- Wasted Count Wasted Size
teness Wasted % Wasted %
Original - 218.3B - 143.9T -
$N=2$ 99.0% 3.8B 1.7% 2.5T 1.7%
$N=3$ 99.2% 7.7B 3.5% 5.1T 3.5%
$N=4$ 99.3% 11.6B 5.3% 7.7T 5.4%
$N=5$ 99.4% 15.4B 7.0% 10.2T 7.1%
Table 1. The post completeness and potentially wasted count and size of posts over the following graph, when only retrieving from the selected NN relays, with N = 2, 3, 4, and 5.
Evaluation of Completeness of Following Edges.
We next evaluate our design from the perspective of users who want to retrieve the post. To look at which posts each retriever should retrieve, we turn to the follower graph. This tells us which posts each user should retrieve. We construct the follower graph $G$, and evaluate the post retrieval completeness over $G$. We evaluate the completeness of each following edge $(U,V)$. Specifically, let $S^{P}_{V}$ denote the set of all posts by $V$, and $O^{P}_{V}$ denote the set of $V$’s posts that can be retrieved by $U$ from the selected $N$ relays for $V$. The completeness is then computed as $({|O^{P}_{V}|}\ /\ {|S^{P}_{V}|})\times 100\%$. Figure 10b shows the CDF of the completeness across the follower edges. Again, we observe that nearly 90% of following edges achieve 100% coverage.
Saved Traffic.  Finally, we evaluate the traffic savings attained by reducing the level of redundant retrievals. To do this, we use the same approach as described in §5.4, while assuming clients only retrieve posts of a user from their selected set of $N$ relays. Specifically, in the original method in §5.4, for a following edge $(U,V)$, we retrieve from $S_{U}^{R}$. This represents the set of relays utilized by user $U$. We then use $S_{V_{N}}^{R}$, which represents the selected $N$ relays for user $V$ (top $N$ relays with largest number of $V$’s posts). Table 1 presents the results for $N=$ 2, 3, 4, and 5. The table shows that even when $N$ is set to 5, it is possible decrease the duplicated retrieval count by 93%, while still achieving 99.4% of post completeness.
Evaluation Summary.  The results highlight that our proposed design, under our simulated scenarios, substantially reduces wasted traffic, without heavily compromising the completeness of the retrieved posts. Although a small (<<1%) number of posts are missed, we posit that this is acceptable in most practical scenarios as the Nostr protocol, by its design, does not guarantee completeness. Considering that numerous relays are lightweight and subject to the costs associated with metered bandwidth, we contend that our proposed design significantly enhances the efficiency and overall sustainability of the Nostr ecosystem, without introducing substantial new complexity.
@@ 7. Related Work
To the best of our knowledge, this research is the first large-scale analysis of the Nostr ecosystem, and Nostr is the first decentralized social networks at scale to prioritize censorship resistance as its primary objective. However, there has been ongoing research on censorship resistance techniques and decentralized social networks.
Censorship Resistance Techniques.  Numerous techniques combat Internet censorship. Proxy-based systems, for instance, use millions of short-lived proxies (Fifield et al., 2012), while TLS-based tools utilize encrypted channels to hinder content monitoring (Frolov and Wustrow, 2019). The Tor network provides anonymity by routing traffic through volunteer-operated servers, concealing user activity (Mani et al., 2018; Wright et al., 2018). Video steganography conceals information within multimedia files for covert communication (Figueira et al., 2022; Barradas et al., 2020). Core network-based circumvention includes methods like decoy routing (Karlin et al., 2011) and domain fronting (Fifield et al., 2015). These varied techniques underscore the multifaceted approach to combating censorship. We complement this prior work with our focus on Nostr, a censorship-resistant microblogging service.
Decentralized Social Networks.  Early efforts in decentralized social networks utilized peer-to-peer (P2P) technologies, with examples like Safebook (Cutillo et al., 2009) and PeerSoN (Buchegger et al., 2009). Recently, federated networks such as Mastodon and PeerTube, forming the Fediverse (Mansoux and Abbing, 2020), have gained prominence, typically using protocols like W3C ActivityPub (W3C, 2018) for server-to-server communication. Research on these federated systems has investigated Mastodon’s centralization tendencies (Raman et al., 2019), growth patterns (Cava et al., 2021), user behaviors (La Cava et al., 2022; La Cava et al., 2022), and content moderation challenges in services like Pleroma (Hassan et al., 2021; Zia et al., 2022). More recently, Bluesky has gained attention, while many of its components are still in a (semi-)centralized mode (Balduf et al., 2025, 2024). Nostr presents a different architecture. Unlike Fediverse applications where servers interoperate, Nostr relays are independent. Users connect directly to multiple relays, enhancing censorship resistance and account portability, as accounts are not tied to a specific relay. This allows users to replicate posts across various relays for better availability.
Other decentralized social networks offer different approaches to this design space. Peer-to-peer (P2P) systems like Secure Scuttlebutt (SSB) (Tarr et al., 2019) use a gossip protocol where users replicate the feeds of those they follow. This shifts the storage burden to the user’s device and can lead to significant client-side overhead and synchronization challenges. Similarly, various blockchain-based social networks (e.g., memo.cash (Zuo et al., 2023), Steemit (Guidi et al., 2021)) embed social data directly onto a blockchain (Arquam et al., 2021; Guidi, 2020; Jiang and Zhang, 2019). While this offers strong immutability, it incurs extreme overhead in terms of storage, transaction fees, and network bandwidth, which must be borne by the entire network. Nostr, while using public-key cryptography for post signing, is distinct: it neither employs blockchain technology (though users can use cryptocurrency for payments), nor complex P2P data replication and validation like SSB. Instead, Nostr prioritizes a simpler, lightweight, and user-friendly design. That said, our findings indicate that Nostr also experiences network overhead, suggesting it is not exempt from the classic resilience-overhead tradeoff inherent in censorship-resistant social networks.
8. Conclusion, Limitations and Future Work
Concluding Remarks.  This paper presents a comprehensive study of the Nostr ecosystem, an emerging censorship-resistant decentralized social network. Nostr represents a new design philosophy for building decentralized social networks: it is intentionally simple and offers minimal features, while prioritizing key attributes that many people expect from a decentralized platform: openness, high availability, and censorship resistance. Overall, we highlight the operational characteristics of such a system in contrast to traditional Fediverse applications, revealing its strengths, limitations, and trade-offs where we also propose potential improvements. We hope these insights can aid in the development of other decentralized social networks, especially for those aiming to prioritize censorship resistance. We wish to flag several limitations, which form the basis of our future work.
Measurement Method Limitations.  The exclusion of offline or paywalled relays from our dataset might skew the results to reflect the behavior of publicly accessible relays. Our findings about relay availability rely on third-party data from nostr.watch. Although this tool is widely used within the Nostr community, conducting independent probes to validate the results would enhance the reliability of our claims regarding relay availability. That said, this approach also offers the advantage of ensuring that the availability results are not influenced by factors such as being banned by relays. The analysis of geolocation and IP-to-AS mapping depends on the IP-API service, which may introduce minor mappuing errors. Thus, future research could improve geolocation accuracy by utilizing multiple sources for geolocation data. Our current estimation of relay operating costs is based on Mastodon instances. To enhance accuracy, future research could improve this by conducting surveys with relay operators. When measuring the availability under attack, we only consider attacks in relay and AS level. However, attacks can occur not only at the AS-level but also at the country level. Studies have examined several instances of country-level censorship, for example, in China (Hoang et al., 2024) and in Turkmenistan (Nourin et al., 2023). Future research may consider these broader levels of attack, which would provide a more comprehensive understanding of availability challenges in different geopolitical contexts.
Proposed Potential Improvements.  Our proposed optimizations are intentionally designed to modify only the client side, a choice made to avoid complicated coordination and ensure easy adoption. A limitation of this client-centric approach is that our evaluation, while demonstrating bandwidth savings, does not measure the potential impact on user-perceived latency or content consistency during query fallbacks. Furthermore, our simulation assumes uniform adoption and does not account for adaptive or alternative user behaviors, such as circumventing replication caps or strategically selecting relay subsets, which could alter real-world performance.
At a higher level, our proposed optimizations, while effective at reducing operational overhead, are not a complete solution to the problem of relay sustainability caused by the lack of incentives. Our work squarely addresses the cost side of the economic equation for operators but does not introduce new revenue streams. Therefore, a critical direction for future work is to explore and evaluate systemic solutions, such as micropayments or tokenization, that provide clear financial incentives to ensure the ecosystem’s long-term health.
Longitudinal Dataset Limitations.  Our dataset captures the Nostr network during a six-month window ending in late 2023. Given Nostr’s rapid evolution, our findings represent a specific point in time, and the network’s characteristics regarding user base, relay stability, and client behavior may have since changed. This highlights critical avenues for future research. A continuous, longitudinal analysis with more recent data is essential to track long-term trends and validate the continued relevance of our conclusions. Further, new decentralized social networks like Bluesky (Balduf et al., 2025, 2024) have emerged recently and attracted a significant number of users. Conducting a comparative performance analysis against these networks would certainly be interesting.
Acknowledgments
This work was supported in part by the Guangzhou Science and Technology Bureau (2024A03J0684), Guangdong provincial project 2023QN10X048, the Guangzhou Municipal Key Laboratory on Future Networked Systems (2024A03J0623), the Guangdong Provincial Key Lab of Integrated Communication, Sensing and Computation for Ubiquitous Internet of Things (No.2023B1212010007), the Guangzhou Municipal Science and Technology Project (2023A03J0011), Guangdong provincial project (2023ZT10X009), and the 111 Center (No. D25008).